OpenStudy (anonymous):

Let f(x) = twenty-four over the quantity of 3 x minus two’ align= . Find f(-2). help!!!

OpenStudy (babalooo):

cool. first step decipher the word problem into a equation of numbers.

OpenStudy (anonymous):

ok... let me plug in what I have, and tell me where I went wrong @babalooo

OpenStudy (anonymous):

f(x)=24/3x-2 y=24/3x-2 x=24/3y-2 what's next???

OpenStudy (babalooo):

quick question, what is this "'align="

OpenStudy (babalooo):

well ignoring the "align" part of the question f(-2) is just asking you to plug in -2 for x of the equation f(x) you don't have to change f(x) into y. just keep the first equation you had posted. and sub in -2 for x and solve for y.

OpenStudy (anonymous):

I'm in an online school, and the lesson told me to replace f(x) with y then switch the y with the x in the equation. have you heard of doing it like that? Because it seems the way you're doing it is different, we plug in numbers at the end when everytihng's switched around, like I came up with -2-22/3=f^-1(-2), initially, it would have been: x-22/3=f^-1(x)

OpenStudy (babalooo):

well, i think your teacher is just getting you guys used to functions. since you are probably more familiar with y=... and letting you guys get more xercise with solving algerbras. I just looked at the equations you had above. the first 2 are correct, but the third is incorrect.

OpenStudy (babalooo):

if y=24/3x then you can just multiply both sides by x/y to get x=24/3y but in the case of y=24/(3x-2), there's a 2 in the bottom so you have to take that into consideration

OpenStudy (anonymous):

I did, and I added 2 to -24 then I got x-22/3=3y/3 then canceled the right side out and got x-22/3=y

OpenStudy (anonymous):

@babalooo

OpenStudy (babalooo):

\[y=\frac{ 24 }{ 3x-2 }\] is this your eqation

OpenStudy (anonymous):

yes

OpenStudy (babalooo):

you see, 3x-2 is one term, you can't split it up like that. secondly, you can't just subtract 2 from a fraction

OpenStudy (babalooo):

\[x-\frac{ 22 }{ 3 } = y\] you got from \[y=\frac{ 24 }{ 3x-2 }\] remember you can only subtract and add fractions when they have the SAME denominator. Here, you have to treat (3x-2) as one term.

OpenStudy (babalooo):

Have you heard of cross multiplying?

OpenStudy (anonymous):

yes, but that isn't the case here, look, I just got off the phone and I'm short on time, if I give you a list of equations, could you do me a favor and finish them by tomorrow?? greatly appreciated

OpenStudy (anonymous):

@babalooo

OpenStudy (babalooo):

lol look, I'm an university student who is doing this for fun in my own spare time. I'm not getting paid for this and I have my own job to work for money. So, if you're looking for someone to do your homework for you at the cost of nothing, where you're not even putting any effort in, I suggest you hire someone to do it and not take advantage of openstudy.

OpenStudy (anonymous):

I understand, sorry for taking advantage of your time, I won't be of further inconvenience to you... and I'm just a Junior in high school